For f(x)=x^2+3x and g(x)=x-2 find f(g(x)).
Answer by stanbon(75887)

For f(x)=x^2+3x and g(x)=x-2 find f(g(x))
---
f[g(x)] = f[x-2] = (x-2)^2 + 3(x-2) = x^2 + 4x + 4 + 6x -12 = x^2+10x-8
=====================
Cheers,
Stan H.
=====================